OpenStudy (moongazer):

Three-fourths of a delivery of fruits were sold on Monday. Half of the remaining unsold were sold on Tuesday. Eight baskets were left and sold on Wednesday. How many basket of fruits were there in all? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

let x baskets after monay..remained=x/4 after tuesday=x/8 x/8=8 x=64

OpenStudy (a_clan):

X - [(3/4)X + (1/2)(X/4)] = 8

OpenStudy (moongazer):

@dipankarstudy please explain

OpenStudy (anonymous):

let x was there ...3/4 was sold..so remained=x(1-3/4)=x/4 half of the remained sold tues....so sold on tuesday=(x/4)[1/2] =x/8 remained=x/4-x/8=x/8 by condition x/8=8 x=64

OpenStudy (moongazer):

@A_clan you are wrong

OpenStudy (a_clan):

i admit , i miscalculated. answer=64

OpenStudy (a_clan):

but method will give the right answer

OpenStudy (moongazer):

thanks dipankarstudy and also A_clan for trying but if you will explain it easier and better i'll also give you a medal :)